Pyridinium bromide (C5H5NHBr) is a strong electrolyte that dissociates completely into C5H5NH+ and Br− . A solution of pyridinium bromide has a pH of 2.88. Kb for pyridine is 1.7×10−9 .What is the molarity of the solution?
Answer #1

pOH = 14 - pH
-log[C5H5NH+ ] = 14 - 2.88
-log[C5H5NH+ ] = 11.12
[C5H5NH+ ] = 10^-11.12
= (7.6*10^-12) M

C5H5NHBr < -- > C5H5NH+ + Br-
c 0 0 intial
c-x x x at equlibrium

x = [C5H5NH+ ]
x = (7.6*10^-12) M
since, x is very small
so, c-x = c
Kb = ([C5H5NH+]*[Br-])/[C5H5NHBr]
1.7*10^-9 = x^2/c
1.7*10^-9 = (7.6*10^-12)^2/c
c = (7.6*10^-12)^2/(1.7*10^-9)
= (3.4*10^-14) M

Answer : (3.4*10^-14) M